# Can Large Language Models Understand Symbolic Graphics Programs?

### v1

@@ 4.3 Benchmarking Semantic Consistency
Figure 6: The semantic consistency test assesses if semantic understanding remains the same when the program is perturbed without semantically changing the image content.
Large language models (LLMs) are exposed to vast amounts of online data, including open-access SVG data. To investigate whether their semantic understanding ability is due to potential data leakage, we propose a semantic consistency test by introducing global translations or rotations to SVG graphics, ensuring SE(2) invariance. Such spatial interventions fundamentally alter the code representation, as SVG graphics consist of lines and Bezier curves with anchor points, and SE(2) operations change all numerical values in the code. However, the SVG’s semantics—such as shape or color—remain unaffected by this perturbation. This allows us to examine how LLMs behave when the same vector graphics are presented with drastic code-numerical changes (see Appendix A.1 for more details). If the model maintains the performance consistency under these perturbations, it may suggest that the semantic understanding is based on a fundamental level of comprehension or visual imagery rather than a trivial memorization of the code.
@@ 4.4 Prediction Entropy of LLMs and Humans
To evaluate the consensus of different LLMs, we compute the average prediction entropy on $500$ symbolic programs using GPT-4o, LLama3-8B, LLama3-70B, Mistral-7B, Yi-1.5-34B, Gemma-1.1-7B and Qwen-1.5-72B. We also conduct a human experiment on the rendered images of these programs and collect the answers (each question has at least 5 participants, see Appendix B). Figure 7 show that humans have strong consensus when answering questions based on images. In comparison, LLMs show low consensus when answering questions based on symbolic programs. This implies that LLMs might have different inner working mechanisms to understand symbolic programs. We are excited by future work to better investigate these differences.
Figure 8: Examples of our instruction data: SVG code, text prompt and paired response (the rendered images are not used for finetuning).
5 Symbolic Instruction Tuning: Boosting Symbolic Program Understanding
Generating symbolic instruction data. Inspired by how visual instruction tuning [54] enables large vision-language models to understand images with visual-question-answering (VQA) data, we design a new method to perform symbolic instruction tuning for LLMs to take steps to better bridge the gap between the semantic understanding and symbolic reasoning over the graphics programs. While to our knowledge there does not exist any semantic instruction-following datasets directly over symbolic programs, these symbolic graphics programs can be rendered into 2D images. With these images in hand, we can easily query powerful vision-language models (e.g., GPT-4o is used in our case) to obtain a detailed semantic captioning based on the rendered image. The intuition is straightforward, as we want to build an intrinsic connection between semantic natural language description and symbolic programs. The instruction data is created in a similar spirit to our benchmark. We leverage the correspondence between symbolic problems and graphics content, and then use the rendered images to obtain semantically rich description. Following this idea, we construct the first semantic description dataset for symbolic graphics programs. Specifically, for each image rendered from a symbolic graphics program, we prompt GPT-4o to produce a semantically-rich description. Finally, we collect a dataset that contains detailed semantic descriptions for 55K symbolic programs.
Supervised finetuning with symbolic instruction data. We generally follow the standard instruction fine-tuning procedure (including the default hyperparameter settings) from Alpaca [90] and use supervised finetuning to train open-source LLMs with our own symbolic instruction data. To facilitate future research, our symbolic instruction data is also made available at Hugging Face.
@@ 5 Symbolic Instruction Tuning: Boosting Symbolic Program Understanding
Experimental results. We perform supervised finetuning on Llama-3.1-8B with orthogonal finetuning [59, 73] to demonstrate the effectiveness of SIT. In Appendix E.2, we provide results on both Llama-3-8B and Gemma-7B to show that the performance gain is agnostic to the base LLM. The performance of LoRA [34] is also given in Appendix E.2 (only slightly worse than OFT). From the experimental results in Table 2, we observe that SIT has consistently improved the semantic graphics program understanding of LLMs, increasing the performance of Llama-3.1-8B from 46.7% to 51.4% with 55K instruction question-answer pairs. With more instruction data (from 10K to 55K), the performance also increases. We note that Llama-3.1-8B achieves competitive performance among all open-source LLMs after being finetuned with our instruction dataset. The performance is already better than GPT-3.5t. The finetuning results demonstrates that the ability of symbolic graphics program understanding can be improved with SIT. However, boosting the performance of open-source LLMs beyond the capability of GPT-4o (the source of our instruction data) remains an open challenge.
6 A Challenge for SVG Program Understanding
Is it really easy to answer semantic reasoning questions over symbolic graphics programs? We provide an intriguing experiment to demonstrate that SVG programs can be quite difficult for LLMs to understand such that even if the corresponding rendered images are fairly easy for humans to recognize, all these powerful LLMs still fail dramatically, only reaching a chance-level accuracy.
@@ 6 A Challenge for SVG Program Understanding
Specifically, we construct symbolic graphics programs that can produce MNIST-like images, as shown in Figure 9. The symbolic programs that can render MNIST-like images are quite simple, typically only containing 1 or 2 path operations. We introduce the SGP-MNIST challenge where we have 100 symbolic graphics programs per digit (in total 10 digits: 0-9). Therefore, we have 1,000 symbolic graphics programs in total. For each problem, we ask one multiple-choice question (with 10 options of digit 0-9): which digit does the SVG program represent? The results are given in Table 3. Despite the fact that these rendered images are quite easy for human to recognize, the corresponding symbolic programs turn out to be difficult for LLMs. Even the powerful GPT-4o can only achieve an accuracy slightly higher than the chance-level. This experiment demonstrates that the kind of symbolic graphics program understanding we explore here can be quite challenging in general, even if some LLMs might seem to work well in some cases. The results suggest that how LLMs understand SVG programs is very different from how humans understand images; better understanding similarities and differences in human and machine reasoning is important if we are to build systems that can appropriately work with us [14]. There are many exciting yet totally unexplored problems in this task, and we hope that our benchmark can serve as a stepping stone to improving the understanding of symbolic graphics problems for LLMs.
7 Improving Generic Instruction Tuning with Symbolic Graphics Programs
Figure 10: Comparison between original and reverse SIT data.
We test the instruction-tuned models on a variety of popular LLM benchmarks, including benchmarks focusing on language understanding (XNLI [15], IFEval [117], HellaSwag [115], C-Eval [37], CoQA [78], MMLU [33], SQuAD2.0 [75]), generic reasoning (BigBenchHard [89], PIQA [6], AGIEval [116]) and mathematical problem-sovling (Arithmetic [8], MathQA [3], GSM8k [12], ASDiv [69]). Specifically, we use the Llama-3.1-8B base model (without any instruction tuning), and then the baseline is finetuned with the Open-Instruct dataset that contains 143K question-answer pairs (details are given in Appendix E.1). We evaluate the effectiveness of using SIT data to improve generic instruction tuning performance by mixing additional SIT data into the Open-Instruct data. Moreover, our SIT data can also be used in a reverse fashion (rev-SIT), i.e., rephrasing the answer as the new question and the question as the new answer. See the comparison between original and reverse SIT data in Figure 10. We test three ways to using SIT data: (1) mixing original SIT data into the Open-Instruct data; (2) mixing the reverse SIT data into the Open-Instruct data; (3) mixing both original and reverse SIT data into the Open-Instruct data. The results are given in Table 4. We can observe that mixing SIT data can generally improve the instruction following ability, suggesting that understanding symbolic graphics programs requires a combination of multiple reasoning abilities and training on SIT data can improve these abilities. More interestingly, we find that the reverse usage of SIT data (i.e., symbolic graphics program generation) can improve a set of reasoning abilities that are complementary to symbolic graphics program understanding. The mixture of both original and reverse SIT data can greatly improve the instruction following ability of LLMs and achieve better performance than the Open-Instruct baseline, the Open-Instruct + SIT baseline and the Open-Instruct + rev-SIT baseline.
Benchmark Open-Instruct Open-Instruct-SIT Open-Instruct-rev-SIT Open-Instruct-mixed-SIT
XNLI [15] 41.8 43.3 (+1.5) 43.1 (+1.3) 42.9 (+1.1)
IFEval [117] 14.8 / 24.9 16.3 (+1.5) / 28.9 (+4.0) 18.3 (+3.5) / 30.5 (+5.6) 16.6 (+1.8) / 29.6 (+4.7)
HellaSwag [115] 60.0 60.2 (+0.2) 60.5 (+0.5) 60.4 (+0.4)
AGIEval [116] 23.7 30.3 (+6.6) 31.6 (+7.9) 29.2 (+5.5)
C-Eval [37] 46.4 47.9 (+1.5) 48.0 (+1.6) 48.1 (+1.7)
BigBenchHard [89] 59.5 60.7 (+1.2) 60.2 (+0.7) 61.2 (+1.7)
Arithmetic [8] 89.8 91.8 (+2.0) 90.1 (+0.3) 91.8 (+2.0)
GSM8k [12] 48.2 50.7 (+2.5) 51.0 (+2.8) 51.5 (+3.3)
CoQA [78] 67.9 69.1 (+1.2) 68.7 (+0.8) 69.1 (+1.2)
PIQA [6] 79.9 80.3 (+0.4) 80.3 (+0.4) 80.4 (+0.5)
MMLU [33] 60.4 61.0 (+0.6) 61.1 (+0.7) 61.6 (+1.2)
MathQA [3] 39.3 40.4 (+1.1) 40.3 (+1.0) 40.7 (+1.4)
SQuAD2.0 [75] 28.9 28.7 (-0.2) 31.6 (+2.7) 29.9 (+1.0)
ASDiv [69] 18.5 21.8 (+3.3) 20.1 (+1.6) 21.3 (+2.8)
Table 4: Results on a variety of popular LLM evaluation benchmarks with or without SIT.
8 Concluding Remarks
@@
We propose a new task over which to evaluate LLMs to contribute towards a multi-faceted assessment of their capabilities: the task of “understand” images – without image; that is, directly from graphics programs. For such a task, we create SGP-Bench, a benchmark that shows distinguishable results between LLMs and introduce the method of Symbolic Instruction Finetuning (SIT) that improves LLMs’ capability of understanding graphics programs. Our work helps deepen our collective understanding of what LLMs can and cannot do, and hopefully inspires a broader diversity of evaluation tasks. However, our work only represents the beginning of symbolic graphics program understanding for LLMs. The inner working mechanisms of such semantic understanding of LLMs and how to improve this ability remain to be explored.
Acknowledgement
The authors would like to thank Yao Feng and Yandong Wen for helpful suggestions. Additionally, HF would like to thank Hanqi Zhou for her support throughout the project, during which time they became engaged (Fun fact: HF and Hanqi Zhou got married a few hours before the submission deadline). The diamond ring featured in Figure 1 symbolizes this joyous personal milestone and is courtesy of the entire team. This work was supported in part by the German Federal Ministry of Education and Research (BMBF): Tubingen AI Center, FKZ: 01IS18039B, and by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. WL was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP XX, project number: 276693517. KMC acknowledges support from the Marshall Scholarship and Cambridge Trust. AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, The Alan Turing Institute, and the Leverhulme Trust via CFI. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH. While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
@@ A.1 Data preparation
Our CAD (3D) sequences data are sampled from DeepCAD [101] datasets, which contains around 180k manually constructed CAD sequences that are originally from the ABC dataset [45]. We manually sample 1000 sequences and use pythonocc (OpenCASCADE) to verify and normalize these CAD sequences, then render the front and back view of the 3D CAD models. Since all the CAD models are from the OnShape platform, the copyright of the CAD models is owned by their creators. For licensing details, see Onshape Terms of Use 1.g.ii. Our CAD (3Drecon.) sequences are sampled from Fusion360 Reconstruction Dataset [97] datasets, which contains 8,625 sequences, with more complex curve operations for constructing sketches. Our CAD (2D) sequences are sampled from SketchGraphs [84] dataset, which consists of 15 million sketches extracted from real-world CAD models. The major difference is it consists of 2D CAD sketches without Extrusion operations.

### v3

@@ 4.3 Benchmarking Semantic Consistency
Figure 10: The semantic consistency test assesses if semantic understanding remains the same when the program is perturbed without semantically changing its rendered content. Image perturbations result in significant code-level changes because symbolic programs rely on absolute coordinates.
Large language models (LLMs) are exposed to vast amounts of online data, including open-access SVG data. To investigate whether their semantic understanding ability is due to potential data leakage, we propose a semantic consistency test by introducing global translations or rotations to SVG graphics, ensuring SE(2) invariance. Such spatial interventions greatly alter the code representation, as SVG graphics consist of lines and Bezier curves with anchor points, and SE(2) operations change all numerical values in the code. However, the SVG’s semantics—such as shape or color—remain unaffected by this perturbation. This allows us to examine how LLMs behave when the same vector graphics are presented with drastic code-numerical changes (see Appendix A.1 for details). Note that we perform non-trivial coordinate-level perturbations to the code, rather than using SVG transformation functions, to prevent shortcut learning by LLMs. Due to the nested structure of the tested SVG code, we visually inspect the perturbed renderings to ensure that the semantics remain unchanged after perturbation. If the model maintains the performance consistency under these perturbations, it suggests that its semantic understanding is based on a fundamental level of comprehension or visual imagery rather than a trivial memorization.
@@ 4.4 Prediction Entropy of LLMs and Humans
To study the consensus of LLMs, we compute the average prediction entropy on $500$ symbolic programs using GPT-4o, LLama3-8B, LLama3-70B, Mistral-7B, Yi-1.5-34B, Gemma-1.1-7B and Qwen-1.5-72B. We conduct a human experiment on the rendered images of these programs and collect the answers (each question has at least 5 participants, see Appendix B). Figure 11 shows that humans have strong consensus when answering questions based on images, while LLMs show low consensus when answering questions based on symbolic programs. This implies that LLMs may have different inner working mechanisms to understand symbolic programs. We are excited by future work to better investigate this difference.
5 Improving LLMS with Symbolic Instruction Tuning
Figure 12: Comparison between original and reverse SIT data.
Generating symbolic instruction data. Inspired by how visual instruction tuning [56] enables large vision-language models to understand images with visual-question-answering (VQA) data, we design a new method to perform symbolic instruction tuning for LLMs to better bridge the gap between the semantic understanding and symbolic reasoning over the graphics programs. To our knowledge, there exist no semantic instruction datasets directly over symbolic graphics programs, After rendering these symbolic graphics programs into images, we can easily query powerful vision-language models (e.g., GPT-4o is used in our case) to obtain a detailed semantic captioning based on the rendered image. The intuition is straightforward, as we want to build an intrinsic connection between semantic description and symbolic programs. The instruction data is created in a similar spirit to our benchmark. We leverage the correspondence between symbolic problems and graphics content, and then use the rendered images to obtain semantically rich description. Following this idea, we construct the first semantic description dataset for symbolic graphics programs. Specifically, for each image rendered from a symbolic graphics program, we prompt GPT-4o to produce a semantically-rich description. Finally, we collect a dataset that contains detailed semantic descriptions for 72K symbolic programs. Moreover, our SIT data can also be used in a reverse fashion (rev-SIT), i.e., rephrasing the answer as the new question and the question as the new answer. Figure 12 shows the comparison between original and reverse SIT data.
Supervised finetuning with symbolic instruction data. We generally follow the standard instruction fine-tuning procedure (including the default hyperparameter settings) from Alpaca [93] and use supervised finetuning to train open-source LLMs with our own symbolic instruction data. To facilitate future research, our symbolic instruction data is also made publicly available.
@@ 5 Improving LLMS with Symbolic Instruction Tuning
Experimental results. We perform supervised finetuning on Llama-3.1-8B with orthogonal finetuning [60, 75] to demonstrate the effectiveness of SIT. Here we use the original SIT data (no rev-SIT data is used). In Appendix E.2, we provide results on both Llama-3-8B and Gemma-7B to show that the performance gain is agnostic to the base LLM. The performance of LoRA [36] is also given in Appendix E.2 (only slightly worse than OFT). From the experimental results in Table 2, we observe that SIT has consistently improved the semantic graphics program understanding of LLMs, increasing the performance of Llama-3.1-8B from 46.7% to 51.4% with 55K instruction question-answer pairs. With more instruction data (from 10K to 55K), the performance also increases. We note that Llama-3.1-8B achieves competitive performance among all open-source LLMs after being finetuned with SIT. The performance is already better than GPT-3.5t. The finetuning results demonstrates that the ability of symbolic graphics program understanding can be improved with SIT. However, the improved performance of Llama-3.1-8B remains worse than Llama-3.1-70B, indicating that the tested ability is fundamental and differences between models of varying scales cannot be easily leveled by finetuning on benchmark-like data.
6 SIT Can Improve General Reasoning Ability
Since we show in Figure 4 that the ability to understand symbolic graphics programs is associated with some fundamental reasoning abilities of LLMs, we are particularly interested in whether symbolic graphcis programs can be used as a novel data source for building better instruction tuning datasets, which can help to improve the general reasoning ability of LLMs. To verify this, we test the instruction-tuned models on a variety of popular LLM benchmarks, including benchmarks focusing on natural language understanding (XNLI [17], IFEval [119], HellaSwag [117], C-Eval [39], CoQA [80], MMLU [34], SQuAD2.0 [77]), generic reasoning (BigBenchHard [92], PIQA [7], AGIEval [118]) and mathematical reasoning (Arithmetic [9], MathQA [3], GSM8k [14], ASDiv [71]).
Benchmark OI OI-SIT OI-rev-SIT OI-mixed-SIT
XNLI 41.8 43.3 (+1.5) 43.1 (+1.3) 42.9 (+1.1)
IFEvalprompt 14.8 16.3 (+1.5) 18.3 (+3.5) 16.6 (+1.8)
IFEvalinst. 24.9 28.9 (+4.0) 30.5 (+5.6) 29.6 (+4.7)
HellaSwag 60.0 60.2 (+0.2) 60.5 (+0.5) 60.4 (+0.4)
C-Eval 46.4 47.9 (+1.5) 48.0 (+1.6) 48.1 (+1.7)
MMLU 60.4 61.0 (+0.6) 61.1 (+0.7) 61.6 (+1.2)
SQuAD2.0 28.9 28.7 (-0.2) 31.6 (+2.7) 29.9 (+1.0)
BBH 59.5 60.7 (+1.2) 60.2 (+0.7) 61.2 (+1.7)
PIQA 79.9 80.3 (+0.4) 80.3 (+0.4) 80.4 (+0.5)
AGIEval 23.7 30.3 (+6.6) 31.6 (+7.9) 29.2 (+5.5)
Arithmetic 89.8 91.8 (+2.0) 90.1 (+0.3) 91.8 (+2.0)
MathQA 39.3 40.4 (+1.1) 40.3 (+1.0) 40.7 (+1.4)
GSM8k 48.2 50.7 (+2.5) 51.0 (+2.8) 51.5 (+3.3)
CoQA 67.9 69.1 (+1.2) 68.7 (+0.8) 69.1 (+1.2)
ASDiv 18.5 21.8 (+3.3) 20.1 (+1.6) 21.3 (+2.8)
Table 3: Results on a variety of popular LLM evaluation benchmarks when performing instruction tuning with or without SIT. The Open-Instruct (OI) dataset serves as our baseline.
Experimental results and discussion. We use the Llama-3.1-8B model (without instruction tuning), and the baseline is finetuned with Open-Instruct (OI) [97] that contains 143K question-answer pairs (details in Appendix E.1). We evaluate whether finetuning with SIT data can improve general reasoning by testing three ways of using SIT data: (1) mixing original SIT data into OI; (2) mixing the reverse SIT data into OI; (3) mixing both original and reverse SIT data into OI. The results are given in Table 3. We can observe that mixing SIT data can generally improve the instruction following and the reverse usage of SIT data (i.e., symbolic graphics program generation) can improve a set of reasoning abilities that are complementary to symbolic graphics program understanding. The mixture of both original and reverse SIT data often achieves better performance than the OI baseline, the OI + SIT baseline and the OI + rev-SIT baseline. These results are consistent with recent findings that training on code can enhance reasoning ability [66, 4] and mathematical understanding [88]. Symbolic graphics programs, a specialized form of code, are used to generate visual graphics content. Like traditional code, they possess a hierarchical structure, but unlike typical programs that produce numerical outputs, symbolic graphics programs generate outputs rich in semantic information, encompassing multiple challenging reasoning tasks such as component localization, color identification, affordance prediction, and semantic and geometric understanding. For instance, answering a question like “What is the object primarily used for?” requires LLMs to first semantically identify the object and then determine its usage. This process involves multiple interconnected reasoning steps, where an error in any one of them leads to an incorrect final answer. SIT enhances reasoning abilities by interpreting low-level graphics programs through high-level natural language descriptions. From Figure 12(a), we see that symbolic graphics program descriptions are highly detailed and semantic—qualities often lacking in general programs.
7 A critical View on current LLM’s Capability
@@
Despite the observed remarkable capability of LLMs to perform complex, multi-step reasoning over symbolic programs, it is evident that there remains substantial potential for further advancements. We provide an intriguing experiment to demonstrate that SVG programs can be quite difficult for LLMs to understand such that even if the corresponding rendered images are fairly easy for humans to recognize, all these powerful LLMs still fail dramatically.
@@
Specifically, we construct symbolic graphics programs that can produce MNIST-like images, as shown in Figure 4 (and Appendix A.1). Our SGP-MNIST dataset contains 1,000 symbolic graphics programs (100 per digit), each asking which digit the SVG program represents. The results are given in Table 4. Even the powerful GPT-4o can only achieve an accuracy slightly higher than the chance-level. The MNIST-like symbolic program presents a unique challenge due to the absence of semantic components for LLMs to reason upon. Instead, it comprises convoluted, irregular path trajectories that resemble handwritten digits. Additionally, the program contains not only single paths but enclosed loops to represent the “thickness” of digits, demanding precise path planning by the LLMs. For instance, the digit 1 is not represented as a single “line” but rather as an elongated loop, which must be distinguished from more oval-shaped loops, such as those representing the digit 0. Without prior knowledge of digit “thickness,” the LLM must infer this distinction through detailed reasoning over the loop structures, further elevating the complexity of the task. The chance-level performance suggests that how LLMs understand SVG programs is very different from how humans understand images; better understanding similarities and differences in human and machine reasoning is important if we are to build systems that can appropriately work with us [16]. There are many exciting yet totally unexplored problems in this task, and our benchmark can serve as a stepping stone to improving symbolic graphics program understanding for LLMs.
8 Related Work and Concluding Remarks
Symbolic graphics programs. Generating visual data by procedural modeling with symbolic programs has been essential to computer graphics since its inception, particularly for 2D shapes and 3D geometry. See [83] for an overview. Common program types include constructive-solid geometry (CSG) [21, 45, 81, 89, 112], computer-aided design (CAD) [30, 50, 51, 87, 108], vector graphics (e.g., SVG) [78, 79], L-systems [33], and customized domains [22, 95, 18, 37, 24, 69]. Among these, SVGs are constructed from primitive shapes like vector paths, curves, or polygons. Central to SVGs is the vector path, providing detailed control over graphics and geometry primitives. Similarly, procedural 3D geometric modeling, particularly in CAD applications, involves parameterized operations to produce geometry. Datasets like the ABC [47] and Fusion 360 Gallery [102] offer hierarchical decomposition, joints, contact surfaces, construction sequences, and shape segmentation based on modeling operations. Our paper focuses on graphics programs of SVG and CAD by introducing a new semantic understanding task that requires a challenging reasoning over the programs.
Relevant benchmarks and datasets. Many benchmarks have evaluated different aspects of LLMs: AI safety/ethics [107, 38], out-of-distribution performance [115, 110], API/tool usage [52], code generation [35], etc. Perhaps the most relevant aspect of LLM evaluation to our task is (non-graphics) program understanding abilities [96, 44, 35, 64, 57, 63, 90, 34, 11]. As graphics programs can be rendered into images, it is also highly relevant to investigate how vision-language models are capable of visual understanding [12, 1, 74, 40, 70, 32, 6, 91, 85, 120]. For SVG programs, [10] studies whether LLMs can understand them and [122] introduces a concurrent benchmark for this purpose. Different from existing benchmarks, SGP-Bench is one of the first benchmarks to evaluate the semantic understanding of general graphics programs.
@@ 8 Related Work and Concluding Remarks
Concluding Remarks. We propose a new task over which to evaluate LLMs to contribute towards a multi-faceted assessment of their capabilities: the task of “understand” images – without image; that is, directly from symbolic graphics programs. For such a task, we create SGP-Bench, a benchmark that shows distinguishable results between LLMs and introduce the method of Symbolic Instruction Finetuning (SIT) that improves LLMs’ capability of understanding graphics programs. More importantly, SIT also helps to improve general reasoning ability of LLMs.
@@ A.1 Data preparation
Our CAD (3D) sequences data are sampled from DeepCAD [105] datasets, which contains around 180k manually constructed CAD sequences that are originally from the ABC dataset [47]. We manually sample 1000 sequences and use pythonocc (OpenCASCADE) to verify and normalize these CAD sequences, then render the front and back view of the 3D CAD models. Since all the CAD models are from the OnShape platform, the copyright of the CAD models is owned by their creators. For licensing details, see Onshape Terms of Use 1.g.ii. Our CAD (3Dcomplex) sequences are sampled from Fusion360 Reconstruction Dataset [101] datasets, which contains 8,625 sequences, with more complex curve operations for constructing sketches. Our CAD (2D) sequences are sampled from SketchGraphs [86] dataset, which consists of 15 million sketches extracted from real-world CAD models. The major difference is it consists of 2D CAD sketches without Extrusion operations.